Solve for x:
(x+3)^2/3=4

Answer by edjones(8007)

(x+3)^2/3=4
3(x+3)^2/3=4*3
(x+3)^2=12
x+3=+-sqrt(12) sqrt of each side.
x+3=+-2sqrt(3)
x=-3+-2sqrt(3)
x= -3+2sqrt(3), x= -3-2sqrt(3)
.
Ed